OpenStudy (anonymous):

write this trinomial in factored form 5a^2+3a-14

OpenStudy (anonymous):

(5a - 7)(a + 2)

OpenStudy (anonymous):

a1 = (-3 + sqrt289)/10 a1 = (-3+17)/10 a1 = 14/10 a2 = -20/10 a2 = -2 The factored form: (a-a1)(a-a2) = (a-14/10)(a + 2)

OpenStudy (anonymous):

(5a - 7)(a + 2)

OpenStudy (anonymous):

5a^2+10a-7a-14 5a(a+2)-7(a+2) (5a-7)(a+2)